# Towards an anomaly detection pipeline for gravitational waves at the Einstein Telescope

### v1

@@ 1 Introduction
The LIGO-Virgo-KAGRA (LVK) Collaboration currently employs both modeled and unmodeled search techniques for gravitational-wave (GW) detection (see LIGOScientific:2025yae for a recent overview), with a preference for modeled searches based on matched filtering. In this approach, precomputed waveform templates representing compact binary coalescences (CBCs) form a template bank that is cross-correlated with the interferometer data. When a template closely matches a potential signal, a trigger is generated, indicating a candidate GW transient for further inspection and validation.
@@
Although highly effective in signal detection, matched filtering is limited by the size and computational cost of the template bank. Since only a limited number of waveform templates can be included in the template bank, mismatches - a measure of how well a template approximates a true signal - are inevitable. It is standard practice to accept mismatches of up to 3% between a real signal and a template (Sakon:2022ibh), corresponding to a reduction in the signal-to-noise ratio (SNR) of up to 1.5%. This seemingly small SNR loss could be critical, especially for signals near the detection threshold, which might be missed.
@@
Unmodeled searches, on the other hand, aim to reconstruct GW signals without relying on specific waveform models. They target a wide variety of sources, and especially short, burst-like signals, which are the most difficult to distinguish from background noise and instrumental artefacts. Such short-duration, hard-to-model signals are ideal targets for machine learning algorithms, which can learn statistical patterns and representations directly from data. Machine and deep learning provide a powerful framework for classification, regression, and inference tasks, and are now widely adopted across most scientific domains. In deep learning, this is typically achieved through multi-layer neural networks trained via gradient-based optimisation to minimise task-specific loss functions (Cuoco:2020ogp; Albertsson:2018maf; Feichtinger:2021uff; Yan:2025hfq; Schafer:2022dxv).
@@
Exploring machine learning approaches is becoming increasingly crucial in light of third-generation GW interferometers such as the Einstein Telescope (ET) (Punturo:2010zz; Hild:2010id; Branchesi:2023mws; Abac:2025saz) or Cosmic Explorer (CE) (Reitze:2019iox; Evans:2021gyd; Gupta:2023lga), representing the European and the United States proposals, respectively, for future ground-based GW detectors. Their improved sensitivity across a broader frequency range will lead to a substantial increase in detectable events, resulting in higher computational demands, not only for simulated data generation but also for detection and analysis of real data. This motivates the development of complementary or alternative search methods to enhance the discovery potential of current and future GW observatories.
@@
We reformulate GW searches as a general anomaly-detection problem: instead of matching data to predefined templates, the algorithm learns the structure of detector noise in frequency–time spectrograms and flags significant deviations as potential signals. In this work, we present the implementation and evaluation of an anomaly-detection algorithm for searching GW transients, based on convolutional autoencoders operating on spectrogram data from a single ET interferometer111As detailed in Sec. 4, for single ET interferometer, we mean that we consider only one out of the 3 detector channels. We focus on burst signals, short GW transients that are produced, for instance, by the mergers of two black holes (BHs) involving or resulting in the formation of an intermediate-mass black hole (IMBH), a class of BHs with masses in the range M⊙ (Greene_2020). We then demonstrate that our model is scalable and applicable across a broader range of BH masses.
@@
The paper is structured as follows. In Sec. 2 we introduce the science case of interest. We detail the network architecture and definition of a simple loss algorithm in Sec. 3. In Sec. 4, we then evaluate the ability of the model to reconstruct ET noise-only spectrograms and define the criteria by which observed data may be flagged as anomalous. We then investigate the performance of the model using the first Einstein Telescope Mock Data Challenge (MDC) (Tania:2025bsa; Regimbau:2012ir) dataset, focusing specifically on BBH mergers where the sum of source frame masses is greater than 100 M⊙. We conclude our discussion in Sec. 5, highlighting future directions of this work.
@@ 2 Intermediate-mass black holes and their gravitational wave signatures
Up to date, the LVK Collaboration has observed 218 GW candidate events, mostly associated with the merger of two black holes, as released in the latest GWTC-4.0 catalog (LIGOScientific:2025slb). Remarkably, several of these mergers involved IMBHs either in the merging components, as in the case of GW190521 (LIGOScientific:2020iuh; LIGOScientific:2020ufj) or in the remnant, as for GW231123 (LIGOScientific:2025rsn).
@@
Second-generation detectors such as LIGO or Virgo are only sensitive to these types of events when the total mass is of the order of up to a few hundred , with the caveat that only the final merger phase can be detected as a ”burst” due to the inspiral phase happening mostly at frequencies below the sensitivities of the detectors. In the case of GW190521 and GW231123 only 4 and 5 cycles were observed in the frequency band of 30-80 Hz, respectively, as reported in (LIGOScientific:2020iuh; LIGOScientific:2020ufj) and in (LIGOScientific:2025rsn); with higher total source-frame masses less cycles will be in the sensitive region of the detector, making these events more difficult to be separated from environmental or instrumental transients (or glitches). ET is designed to increase the sensitivity to GW strains of at least an order of magnitude with respect to second-generation detectors, and will therefore be able to capture GW signals happening at lower frequencies, also thanks to its cryogenic operations, which in turn will allow us to extend to higher total masses the sensitivity of the detector. In ET, events like GW190521 or GW231123 will be detected with the final phase of the inspiral, typically staying within the detector-sensitive region for up to a few seconds. For events with total masses up to 0.5-1 M⊙, however, only a few cycles of the final merger phase will be detected by ET, and more massive events will fall outside the sensitivity of ET. Interestingly, space-based observatories planned to be operating at the same time of ET or CE, as LISA, will be able to capture the inspiral phase of such massive mergers, with the possibility of multi-band detections (Abac:2025saz).
@@ 3 Autoencoders as anomaly detection tools
Autoencoders (AEs) are a class of unsupervised learning algorithms in which input data are encoded into a generally low-dimensional latent space and subsequently decoded back into the original data domain (2020arXiv200305991B). Through this process, the AE learns a compressed internal representation that captures the most essential features of the input. The quality of the reconstruction after encoding and decoding is quantified by the reconstruction error, typically measured using the mean-squared-error (MSE) loss function:
@@
Several variants of autoencoders have been developed to adapt the basic architecture to different types of data and tasks. Convolutional autoencoders (CAEs) are well-suited for structured inputs such as images or spectrograms, as they employ convolutional layers to extract spatially localised features (Morawski:2021kxv; Pol:2018nhq; Gandrakota:2024yqs; Bacon:2022lsm). Variational autoencoders (VAEs) (2019arXiv190602691K; Oshino:2025laz) introduce a probabilistic framework by learning a distribution over the latent space, enabling generative modelling and smooth interpolation between data points. Other extensions include denoising autoencoders (DAEs) (2013arXiv1305.6663B), which learn to reconstruct clean inputs from noisy versions, and sparse autoencoders (SAEs) (2023arXiv230908600C), which promote sparsity in the latent representation.
The choice of autoencoder type depends on the structure of the input data and the intended application. In our case, we adopt CAEs because our input data consists of images, specifically spectrograms (time vs. frequency), which contain noise. We opt for a minimal architecture to minimise computational cost and training time, as shown in Fig. 2. We implemented a deep CAE in PyTorch (Paszke:2019xhz) to learn a compact representation of spectrogram images and perform anomaly detection. The model consists of three convolutional encoding blocks followed by three symmetric decoding blocks; each encoding block comprises a convolutional layer (with kernel size 3 and padding 1), batch normalisation, dropout to prevent the model from focusing on some specific features or neurons overfitting the data, and max pooling (with indices retained for unpooling). The decoder mirrors the encoder using max unpooling layers, followed by transposed convolutions, batch normalisation, and dropout, to reconstruct the input.
@@ 3.1 Introduction of weak supervision
We introduce the use of Optuna in optimisation of the model hyperparameters (akiba2019optuna), including the newly defined separation margin, . A value of was found to induce strong separation between noise-only and signal-injected spectrograms, while maintaining stability of the training process. Higher values of were found to introduce instability into the results. We notice that in other fields, such as high-energy physics or collider phenomenology, a similar approach to optimising losses but with different function choices was implemented for example in Khosa:2020qrz; Banda:2025nrv; Park:2020pak.
@@ 4.1 Noise modelling
This study utilises the Einstein Telescope mock data challenge dataset, which corresponds to approximately one month of simulated observations (Tania:2025bsa). This dataset is provided as a time series sampled at 8192 Hz, segmented into 2048-second chunks, and includes realistic noise strains for the three interferometers , , and in ET’s triangular configuration. Since the final location of the detector is not known yet, the mock data were prepared by positioning ET on the Virgo site in Cascina, with the vertex coincident with the vertex of Virgo as described in (Tania:2025bsa; Regimbau:2012ir).
For simplicity, this study does not yet implement the antenna patterns of the detector and restricts analysis to only the polarisation component. We limit ourselves to selecting only one detector channel (E1), noting that the choice among any of the three detectors has a negligible impact. We estimate the power spectral density (PSD) using Welch’s method (an averaged FFT-based estimator) and whitened the data accordingly. The whitened time series is then split into 2-second segments to construct spectrograms. An example of such is shown in fig. 3.
@@ 4.2 Signal injection
Having developed an understanding of how the model handles noise-only spectrograms, we now estimate its ability to distinguish anomalous input data. The ET MDC data set comprises a set of GW signals including 59,540 binary neutron star (BNS), 6,578 binary black hole (BBH), and 1,977 neutron star–black hole (NSBH) mergers. Here, we focus solely on BBH mergers, specifically those involving or forming IMBHs. The MDC data also provides spin, eccentricity and inclination parameters. We use the IMRPhenomPV2 approximant and include these in the generation of signal waveforms. Figure 6 provides an example of the signal injection process, with spectrograms depicting the signal only (top), noise only (bottom), and the full signal-injected example (middle).
@@ 5 Conclusions and Outlook
This last point, however, opens the way to further optimisation. In the current design, the algorithm is applied to only one detector; however, it can also be applied to several detectors in a network, thereby reducing the possibility of statistical fluctuations in the noise or glitches if the detectors are not co-located. Furthermore, it would be possible to introduce an additional module to classify anomalies as possible signals, as done, for example, in (Bini:2023gil), and eventually also estimate their parameters (Dax:2021tsq; Papalini:2025exy; Lanchares:2025kpb). A combination of multiple machine learning modules aiming at anomaly detection of signal candidates, signal candidates classification, and parameter estimation, in a single framework, will provide a fully automated analysis pipeline capable of detecting GWs and generating low-latency alerts.
@@ Appendix A Ringdown frequency
To relate the merger (ringdown) frequency to the total source–frame mass of the coalescing binary, we adopt the expression for the dominant quadrupolar quasinormal mode (QNM) of the final black hole (Berti:2005ys). The frequency of the mode can be written as:
@@
Assuming a typical final spin of , we obtain (Keitel:2016krm; Healy:2014eua). The remnant mass is then estimated from the source–frame total mass by accounting for the radiated energy and cosmological redshift as
with representing the fraction of mass–energy lost in the form of gravitational waves (Keitel:2016krm; Healy:2014eua). Combining these expressions yields the working formula:
@@ Appendix B Implementation details
The initial development of the CAE model described herein was conducted using simple 2-second-long white noise spectrograms generated with standard deviation of . A total of spectrograms are generated and split 70%/30% for training/testing purposes. BBH merger signals are generated with the IMRPhenomHM approximant, with only masses and distance considered for simplicity. Figure 11 shows spectrograms containing white-noise-only (bottom), generated signal (top) and this signal injected into the noise (middle).

### v2

@@ 1 Introduction
The LIGO-Virgo-KAGRA (LVK) Collaboration currently employs both modeled and unmodeled search techniques for gravitational-wave (GW) detection (see Abac and others (2025b) for a recent overview), with a preference for modeled searches based on matched filtering. In this approach, precomputed waveform templates representing compact binary coalescences (CBCs) form a template bank that is cross-correlated with the interferometer data. When a template closely matches a potential signal, a trigger is generated, indicating a candidate GW transient for further inspection and validation.
@@
Although highly effective in signal detection, matched filtering is limited by the size and computational cost of the template bank. Since only a limited number of waveform templates can be included in the template bank, mismatches - a measure of how well a template approximates a true signal - are inevitable. It is standard practice to accept mismatches of up to 3% between a real signal and a template (Sakon and others, 2024), corresponding to a reduction in the signal-to-noise ratio (SNR) of up to 1.5%. This seemingly small SNR loss could be critical, especially for signals near the detection threshold, which might be missed.
@@
Unmodeled searches, on the other hand, aim to reconstruct GW signals without relying on specific waveform models. They target a wide variety of sources, and especially short, burst-like signals, which are the most difficult to distinguish from background noise and instrumental artefacts. Such short-duration, hard-to-model signals are ideal targets for machine learning algorithms, which can learn statistical patterns and representations directly from data. Machine and deep learning provide a powerful framework for classification, regression, and inference tasks, and are now widely adopted across most scientific domains. In deep learning, this is typically achieved through multi-layer neural networks trained via gradient-based optimisation to minimise task-specific loss functions (Cuoco and others, 2021; Albertsson and others, 2018; Feichtinger and others, 2022; Yan et al., 2025; Schäfer and others, 2023).
@@
Exploring machine learning approaches is becoming increasingly crucial in light of third-generation GW interferometers such as the Einstein Telescope (ET) (Punturo and others, 2010; Hild and others, 2011; Branchesi and others, 2023; Abac and others, 2025d) or Cosmic Explorer (CE) (Reitze and others, 2019; Evans and others, 2021; Gupta and others, 2024), representing the European and the United States proposals, respectively, for future ground-based GW detectors. Their improved sensitivity across a broader frequency range will lead to a substantial increase in detectable events, resulting in higher computational demands, not only for simulated data generation but also for detection and analysis of real data. This motivates the development of complementary or alternative search methods to enhance the discovery potential of current and future GW observatories.
@@
We reformulate GW searches as a general anomaly-detection problem: instead of matching data to predefined templates, the algorithm learns the structure of detector noise in frequency–time spectrograms and flags significant deviations as potential signals. In this work, we present the implementation and evaluation of an anomaly-detection algorithm for searching GW transients, based on convolutional autoencoders operating on spectrogram data from a single ET interferometer111As detailed in Sec. 4, for single ET interferometer, we mean that we consider only one out of the 3 detector channels. We focus on burst signals, short GW transients that are produced, for instance, by the mergers of two black holes (BHs) involving or resulting in the formation of an intermediate-mass black hole (IMBH), a class of BHs with masses in the range M⊙ (Greene et al., 2020). We then demonstrate that our model is scalable and applicable across a broader range of BH masses.
@@
The paper is structured as follows. In Sec. 2 we introduce the science case of interest. We detail the network architecture and definition of a simple loss algorithm in Sec. 3. In Sec. 4, we then evaluate the ability of the model to reconstruct ET noise-only spectrograms and define the criteria by which observed data may be flagged as anomalous. We then investigate the performance of the model using the first Einstein Telescope Mock Data Challenge (MDC) (Regimbau and Suresh, 2025; Regimbau and others, 2012) dataset, focusing specifically on BBH mergers where the sum of source frame masses is greater than 100 M⊙. We conclude our discussion in Sec. 5, highlighting future directions of this work.
@@ 2 Intermediate-mass black holes and their gravitational wave signatures
Up to date, the LVK Collaboration has observed 218 GW candidate events, mostly associated with the merger of two black holes, as released in the latest GWTC-4.0 catalog (Abac and others, 2025c). Remarkably, several of these mergers involved IMBHs either in the merging components, as in the case of GW190521 (Abbott and others, 2020b, c) or in the remnant, as for GW231123 (Abac and others, 2025a).
@@
Second-generation detectors such as LIGO or Virgo are only sensitive to these types of events when the total mass is of the order of up to a few hundred , with the caveat that only the final merger phase can be detected as a ”burst” due to the inspiral phase happening mostly at frequencies below the sensitivities of the detectors. In the case of GW190521 and GW231123 only 4 and 5 cycles were observed in the frequency band of 30-80 Hz, respectively, as reported in (Abbott and others, 2020b, c) and in (Abac and others, 2025a); with higher total source-frame masses less cycles will be in the sensitive region of the detector, making these events more difficult to be separated from environmental or instrumental transients (or glitches). ET is designed to increase the sensitivity to GW strains of at least an order of magnitude with respect to second-generation detectors, and will therefore be able to capture GW signals happening at lower frequencies, also thanks to its cryogenic operations, which in turn will allow us to extend to higher total masses the sensitivity of the detector. In ET, events like GW190521 or GW231123 will be detected with the final phase of the inspiral, typically staying within the detector-sensitive region for up to a few seconds. For events with total masses up to 0.5-1 M⊙, however, only a few cycles of the final merger phase will be detected by ET, and more massive events will fall outside the sensitivity of ET. Interestingly, space-based observatories planned to be operating at the same time of ET or CE, as LISA, will be able to capture the inspiral phase of such massive mergers, with the possibility of multi-band detections (Abac and others, 2025d).
@@ 3 Autoencoders as anomaly detection tools
Autoencoders (AEs) are a class of unsupervised learning algorithms in which input data are encoded into a generally low-dimensional latent space and subsequently decoded back into the original data domain (Bank et al., 2020). Through this process, the AE learns a compressed internal representation that captures the most essential features of the input. The quality of the reconstruction after encoding and decoding is quantified by the reconstruction error, typically measured using the mean-squared-error (MSE) loss function:
@@
Several variants of autoencoders have been developed to adapt the basic architecture to different types of data and tasks. Convolutional autoencoders (CAEs) are well-suited for structured inputs such as images or spectrograms, as they employ convolutional layers to extract spatially localised features (Morawski et al., 2021; Pol et al., 2019; Gandrakota, 2025; Bacon et al., 2023). Variational autoencoders (VAEs) (Kingma and Welling, 2019; Oshino et al., 2025) introduce a probabilistic framework by learning a distribution over the latent space, enabling generative modelling and smooth interpolation between data points. Other extensions include denoising autoencoders (DAEs) (Bengio et al., 2013), which learn to reconstruct clean inputs from noisy versions, and sparse autoencoders (SAEs) (Cunningham et al., 2023), which promote sparsity in the latent representation.
The choice of autoencoder type depends on the structure of the input data and the intended application. In our case, we adopt CAEs because our input data consists of images, specifically spectrograms (time vs. frequency), which contain noise. We opt for a minimal architecture to minimise computational cost and training time, as shown in Fig. 2. We implemented a deep CAE in PyTorch (Paszke and others, 2019) to learn a compact representation of spectrogram images and perform anomaly detection. The model consists of three convolutional encoding blocks followed by three symmetric decoding blocks; each encoding block comprises a convolutional layer (with kernel size 3 and padding 1), batch normalisation, dropout to prevent the model from focusing on some specific features or neurons overfitting the data, and max pooling (with indices retained for unpooling). The decoder mirrors the encoder using max unpooling layers, followed by transposed convolutions, batch normalisation, and dropout, to reconstruct the input.
@@ 3.1 Introduction of weak supervision
We introduce the use of Optuna in optimisation of the model hyperparameters (Akiba et al., 2019), including the newly defined separation margin, . A value of was found to induce strong separation between noise-only and signal-injected spectrograms, while maintaining stability of the training process. Higher values of were found to introduce instability into the results. We notice that in other fields, such as high-energy physics or collider phenomenology, a similar approach to optimising losses but with different function choices was implemented for example in Khosa and Sanz (2023); Banda et al. (2025); Park et al. (2020).
@@ 4.1 Noise modelling
This study utilises the Einstein Telescope mock data challenge dataset, which corresponds to approximately one month of simulated observations (Regimbau and Suresh, 2025). This dataset is provided as a time series sampled at 8192 Hz, segmented into 2048-second chunks, and includes realistic noise strains for the three interferometers , , and in ET’s triangular configuration. Since the final location of the detector is not known yet, the mock data were prepared by positioning ET on the Virgo site in Cascina, with the vertex coincident with the vertex of Virgo as described in (Regimbau and Suresh, 2025; Regimbau and others, 2012).
For simplicity, this study does not yet implement the antenna patterns of the detector and restricts analysis to only the polarisation component. We limit ourselves to selecting only one detector channel (E1), noting that the choice among any of the three detectors has a negligible impact. We estimate the power spectral density (PSD) using Welch’s method (an averaged FFT-based estimator) and whiten the data accordingly. The whitened time series is then split into 2-second segments to construct spectrograms. An example of such is shown in fig. 3.
@@ 4.2 Signal injection
Having developed an understanding of how the model handles noise-only spectrograms, we now estimate its ability to distinguish anomalous input data. The ET MDC data set comprises a set of GW signals including 59,540 binary neutron star (BNS), 6,578 binary black hole (BBH), and 1,977 neutron star–black hole (NSBH) mergers. Here, we focus solely on BBH mergers, specifically those involving or forming IMBHs. The MDC data also provides spin, eccentricity and inclination parameters. We use the IMRPhenomPv2 approximant and include these in the generation of signal waveforms (Khan et al., 2016). For signal-injected samples, the merger time is required to lie within the spectrogram window, but its position is randomly shifted within the segment (up to 30% around the center of the window) in order to prevent the model from learning a fixed time alignment and to reduce overfitting. Figure 6 provides an example of the signal injection process, with spectrograms depicting the signal only (top, illustratively placed in the center of the spectrogram), noise only (bottom), and the full signal-injected example (middle).
@@ 5 Conclusions and Outlook
This last point, however, opens the way to further optimisation. In the current design, the algorithm is applied to only one detector; however, it can also be applied to several detectors in a network, thereby reducing the possibility of statistical fluctuations in the noise or glitches if the detectors are not co-located. Furthermore, it would be possible to introduce an additional module to classify anomalies as possible signals, as done, for example, in (Bini et al., 2023), and eventually also estimate their parameters (Dax et al., 2021; Papalini et al., 2025; Lanchares et al., 2025). A combination of multiple machine learning modules aiming at anomaly detection of signal candidates, signal candidates classification, and parameter estimation, in a single framework, will provide a fully automated analysis pipeline capable of detecting GWs and generating low-latency alerts.
@@ Appendix A Ringdown frequency
To relate the merger (ringdown) frequency to the total source–frame mass of the coalescing binary, we adopt the expression for the dominant quadrupolar quasinormal mode (QNM) of the final black hole (Berti et al., 2006). The frequency of the mode can be written as:
@@
Assuming a typical final spin of , we obtain (Keitel and others, 2017; Healy et al., 2014). The remnant mass is then estimated from the source–frame total mass by accounting for the radiated energy and cosmological redshift as
with representing the fraction of mass–energy lost in the form of gravitational waves (Keitel and others, 2017; Healy et al., 2014). Combining these expressions yields the working formula:
@@ Appendix B Implementation details
The initial development of the CAE model described herein was conducted using simple 2-second-long Gaussian-like white noise spectrograms generated with standard deviation of , using a standard short-time Fourier transform. A total of spectrograms are generated and split 70%/30% for training/testing purposes. BBH merger signals are generated with the IMRPhenomPv2 waveform approximant (Khan et al., 2016), with only masses and distance considered as parameters, for simplicity. For signal-injected samples, the merger time is required to lie within the spectrogram window, but its position is randomly shifted within the segment (up to 30% around the center of the window) in order to prevent the model from learning a fixed time alignment and to reduce overfitting. No additional optimisation of the time–frequency representation was performed.
Although the IMBH signals studied herein inherently inhabit only the lowest frequency bins of the spectrograms, we elect not to apply any form of low-pass filters or cuts. This is done because, while this study specifically targets IMBH mergers, avoiding any such early selection criteria preserves flexibility for any future expansion of the method to a wider range of signal morphologies. Figure 11 shows spectrograms containing white-noise-only (bottom), generated signal (top, illustratively placed in the center of the spectrogram) and this signal injected into the noise (middle).